OpenStudy (anonymous):

I need help solving for x using complete the square 3x^2-18x=12 @precal

OpenStudy (anonymous):

Step 1: Divide by 3: x^2 - 6x = 4 Step 2: Make the Square: (x-3)^2 = 4 This gets you the square part but it isn't equal to step 1. Step 3: Make the equation equivalent (x-3)^2 = x^2 - 6x + 9 so the leftover part is 9. So to do this you just add 9 to the other side to get 13 (x-3)^2 = 13 Step 4: Take the Square Root of both sides x-3 = + or - √13 Step 5: Make x on one side: x = 3 + √13 AND x = 3 - √13 That's about it.

OpenStudy (precal):

can't ignore the 3 in the front

OpenStudy (precal):

I will look at this later, have many things to do ........sorry deadlines.......

OpenStudy (precal):